OpenStudy (anonymous):

2 cubic yards = ? cubic inches?

OpenStudy (anonymous):

1 (cubic yard) = 46.656 cubic inches so you can take it from there

OpenStudy (lgbasallote):

\[2 \; yd^3 \times \frac{3 ft}{1 yd} \times \frac{3 ft}{1 yd} \times \frac{ 3 ft}{1 yd}\]

OpenStudy (anonymous):

okay thanks.

OpenStudy (lgbasallote):

just simplify that

OpenStudy (lgbasallote):

then divide by 12^3

OpenStudy (anonymous):

@lgbasallote i think you mean multiply

OpenStudy (lgbasallote):

no wait..times

OpenStudy (lgbasallote):

sorry

OpenStudy (anonymous):

the final answer is 93312

OpenStudy (agent47):

you should read up on how to convert units

OpenStudy (lgbasallote):

if you're not interested to see how to do it, google does it fas

OpenStudy (lgbasallote):

fast*

OpenStudy (anonymous):

thanks.

hero (hero):

Alternative way to do it:

hero (hero):

\(2 {yd}^3 = x \dot\ {in}^3\) \[\frac{2}{x} = \frac{in^3}{yd^3}\] \[\frac{2}{x} = (\frac{in}{yd})^3\] \[\sqrt[3]{\frac{2}{x}} = \frac{in}{yd}\] \[\sqrt[3]{\frac{2}{x}} = \frac{in}{36in}\] \[\sqrt[3]{\frac{2}{x}} = \frac{1}{36}\] \[{\frac{2}{x}} = (\frac{1}{36})^3\] \[{\frac{2}{x}} = \frac{1}{46656}\] x = 93312

hero (hero):

@RuSSian1909